OpenStudy (idealist10):

Solve y'=(-6x+y-3)/(2x-y-1) implicitly.

OpenStudy (idealist10):

@ganeshie8

OpenStudy (idealist10):

y=ux u=y/x y'=u'x+u u'x+u=(-6x+ux-3)/(2x-ux-1)=((x(u-6))-3)/((x(2-u))-1)

ganeshie8 (ganeshie8):

you need to have DE in homogeneous form before doing that

ganeshie8 (ganeshie8):

Is the given DE homogeneous ?

OpenStudy (idealist10):

It seems no.

ganeshie8 (ganeshie8):

why ?

OpenStudy (idealist10):

Because I just did the substitution but it didn't work.

ganeshie8 (ganeshie8):

did u get that method ?